Analysis of a rock sample shows that only 1/16 of the original amount of chlorine-36 remains in
the rock. Estimate the age of the rock given that the half-life of chlorine-36 is 3.0 x 105
a
1
Expert's answer
2020-06-04T09:58:39-0400

The number of atoms N remaining after time t is given by the law od radioactive decay:


"N = N_02^{-\frac{t}{T}}"

where "N_0" is the number of atoms at the initial moment of time, "T" is the half-life.

Expressing "t" (age) from this equation:


"t = T\log_2\dfrac{N_0}{N}"

As far as "N\/N_0 = 1\/16" and "T = 3\cdot 10^{5} years", obtain:


"t = 3\cdot 10^{5}\log_216 = 3\cdot 10^{5}\cdot 4 = 1.2\cdot 10^6years"

Answer. The age of the sample is 1.2*10^6 years.